# BOWIE-ALIGN: A JWST comparative survey of aligned vs misaligned hot Jupiters to test the dependence of atmospheric composition on migration history–A BOWIE-ALIGN: A JWST comparative survey of aligned vs misaligned hot Jupiters to test the dependence of atmospheric composition on migration history

### v1

@@ Abstract
A primary objective of exoplanet atmosphere characterisation is to learn about planet formation and evolution, however, this is challenged by degeneracies. To determine whether differences in atmospheric composition can be reliably traced to differences in evolution, we are undertaking a new survey with JWST to compare the compositions of a sample of hot Jupiters that orbit F stars above the Kraft break with different orbital alignments. Under the assumption that aligned planets migrate through the inner disc, while misaligned planets migrate after disc dispersal, the act of migrating through the inner disc should lead to a measurable difference in the C/O between aligned and misaligned planets. We expect the amplitude and sign of this difference to depend on the amount of planetesimal accretion and whether silicates accreted from the inner disc release their oxygen. Here, we identify all known exoplanets that are suitable for testing this hypothesis, describe our JWST survey, and use noise simulations and atmospheric retrievals to estimate our survey’s sensitivity. With the selected sample of four aligned and four misaligned hot Jupiters, we will be sensitive to the predicted differences in C/O between aligned and misaligned hot Jupiters for a wide range of model scenarios.
@@
keyword1 – keyword2 – keyword31 Introduction
@@
Penzlin & Booth et al. 2024 used simulations to demonstrate that the unconstrained values of key disc and planet formation parameters such as dust-to-gas mass, disc temperature, and the relative drift of dust to gas (Stokes-to-alpha number), create a degeneracy between C/O and metallicity (defined here as O/H) over a wide dynamic range. These uncertainties are hard to constrain robustly through independent observations. Therefore, we propose the best way to determine whether differences in formation history lead to a measurable difference in atmospheric C/O and metallicity is by comparing populations of planets for which we have independent evidence that they underwent different evolutionary pathways. Specifically, as Penzlin & Booth et al. 2024 show, planets that have undergone disc-free (high eccentricity) migration should have different C/O and metallicity to planets that have undergone disc migration. The idea behind this is that disc-migrated planets will accrete solids from the inner disc during their migration while disc-free, high-eccentricity migrated planets will not since they complete their migration after disc dispersal (Figure 1).
@@ 2 Target selection
To test whether atmospheric composition depends on migration history, we needed to construct a sample of planets with large predicted atmospheric signals and precisely measured obliquities. To do this, we drew from the TEPCAT catalogue111https://www.astro.keele.ac.uk/jkt/tepcat/obliquity.html, accessed 23rd May 2024 (Southworth, 2011) of 180 exoplanets with measured sky-projected obliquities, , primarily measured via the Rossiter-McLaughlin effect (Rossiter, 1924; McLaughlin, 1924). We define misaligned exoplanets as those with sky-projected obliquities ∘ and aligned planets as those with ∘ and only considered those planets with precisely measured obliquities (() ∘). While we based our sample selection on sky-projected obliquities, it is possible that planets with low sky-projected obliquities could, in fact, be misaligned in 3D space once the stellar inclination is resolved. However, of the 20 planets with measured sky-projected and 3D obliquities from the TEPCAT catalogue (Southworth, 2011), only one (5 % of the total) is aligned in 2D, sky-projected space and is misaligned in 3D space. Therefore, our target selection based on sky-projected obliquities is unlikely to be biased by potential 3D misalignments.
@@
Of these planets, we are acquiring new JWST observations of TrES-4b, KELT-7b, NGTS-2b, HAT-P-30b and WASP-15b through Programme GO 3838 (PIs: Kirk & Ahrer) with the other targets already observed or due to be observed by other approved programs (WASP-17b, GTO 1353, PI: Lewis; HD 149026b, GTO 1274, PI: Lunine, Bean et al. 2023; WASP-94Ab, GO 3154, PI: Ahrer). All of these planets have eccentricities consistent with zero. With this combination of archival and new data, we will obtain (i) a four planet sample of aligned planets that likely migrated through the disc and (ii) a four planet sample of misaligned planets that ended their formation at orbits near and beyond the water ice line, and reach their final orbits after disc dispersal.
@@ 3 Observing strategy
We will observe our targets in transmission using the JWST NIRSpec G395H instrument mode, which covers several absorption bands from the primary O and C bearing species in hot Jupiters (H2O, CO2 and CO, Figure 3 and e.g., Alderson et al. 2023). This approach will allow us to measure C/O and O/H and thus address whether these quantities differ between aligned and misaligned hot Jupiters. While NIRSpec/PRISM would cover the same molecular features in addition to covering bluer wavelengths (albeit at lower spectral resolution), it saturates more quickly than NIRSpec/G395H. The JWST Early Release Science observations of WASP-39b (K mag = 10.2) with NIRSpec/PRISM saturated at near-IR wavelengths (Rustamkulov et al., 2023), with the data synthesis analysis recommending the avoidance of partial saturation with PRISM (Carter & May et al., in press). Since four of our five GO 3838 targets are brighter than WASP-39 (Table 1), we selected G395H to avoid saturation and to ensure a homogeneous data set across the sample. Similarly, the observations of WASP-17b and WASP-94Ab also make use of the G395H instrument. However, HD 149026b was observed with NIRCam in emission (Bean et al., 2023), and therefore, we will need to take this into account when we combine the full sample. We discuss the results of Bean et al. (2023) in Section 6.
@@ 4 Predictions from disc models and 1D atmosphere models
For our actual sample size of eight planets, we estimate a precision on C/O of (aligned planets) and (misaligned planets) for the solar metallicity, fiducial case (disc scenario 1). This would make the predicted difference between the aligned and misaligned planets of particularly challenging to measure. However, if the planets are more metal-rich, similar to other planets in the literature, then we would be sensitive to the difference between the C/O of aligned and misaligned planets in the solar metallicity fiducial setup. Specifically, we estimate a precision on C/O of (aligned planets) and (misaligned planets) in this case, which would allow us to measure the predicted difference between the C/O of the two populations of at .
@@ 7 Additional science enabled by our homogeneous sample
In addition to the 1D forward and retrieval modelling we will perform for our sample, we will investigate the atmospheres of our exoplanets in 3D. Since all of our targets are on short orbital periods they are expected to be tidally locked. This leads to a large day-to-night temperature gradient between their permanent daysides and permanent nightsides which, in turn, drives a super-rotating equatorial jet that carries hot dayside gas to the nightside and cool nightside gas to the dayside (e.g., Showman & Polvani, 2010). The result is hotter evening limbs and cooler morning limbs, and hence different gas phase chemistry and cloud coverage at each limb. Using a combination of ingress/egress light curve fits and light curve fits with two planetary radii for the morning and evening limb, Espinoza et al. (in press) have measured limb differences in a planet’s transmission spectrum with JWST. We will apply the same techniques, using Tiberius and catwoman101010https://github.com/KathrynJones1/catwoman (Jones & Espinoza, 2020) to search for limb asymmetries in our data.
@@
Our investigations of 3D effects and photochemistry will be backed up by a comparison of each planet’s observed spectrum to the spectrum predicted by a 3D climate model (also known as a general circulation model, or GCM) of a hot Jupiter atmosphere. This comparison will reveal how the interactions between atmospheric circulation, radiative transfer, disequilibrium thermochemistry (e.g., Zamyatina et al., 2024) and photochemistry, in 3D, impact the exoplanets’ morning and evening limb spectra. Furthermore, this will allow us to determine to what extent the limb-averaged C/O and Z we measure from our transit spectra are representative of the overall atmosphere composition. Among the GCM outputs will be longitude-latitude maps of molecular abundances as a function of pressure and temperature, which we will use to inform our 1D models. The ALIGN programme sample could also serve as a high-quality observational reference for model intercomparison projects, such as those within the CUISINES framework (Fauchez et al., 2021; Sohl et al., 2024). The homogeneity of data would allow to benchmark single-column models and 3D GCMs in terms of e.g. their simulated transmission spectra, for aligned and misaligned hot Jupiters. This will help to validate our theoretical predictions and make the conclusions of the programme more robust.
@@ 8 Summary
: this result would indicate that the majority of hot Jupiter assembly is completed before migration.
In this paper, we have described our observational test of these predictions, including the target selection and sample size, which we investigated using a combination of 256 1D atmospheric forward models, JWST noise simulations and atmospheric retrievals. Our simulated transmission spectra and retrievals show that we will be sensitive to result (i) if the atmospheric metallicities approach solar, while the greater differences in C/O between the two populations for result (ii) means we will be sensitive to these over a wide range of metallicities. If we measure either result (i) or (ii), this would represent the first observational confirmation that atmospheric composition traces formation environment. If we observe super-solar metallicities, this would indicate the accretion of solid material, like the solar system, while sub-solar metallicities would suggest that hot Jupiter assembly differs from the giant planets in the solar system.
Our comparative survey of a population of carefully-selected hot Jupiters will robustly test the dependence of atmospheric composition on formation in a non-degenerate way.
@@ Acknowledgements
J.K. acknowledges financial support from Imperial College London through an Imperial College Research Fellowship grant. J.E.O. and R.A.B are supported by Royal Society University Research Fellowships and A.B.T.P. received support from their Enhancement Awards. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Framework Programme (grant agreement no. 853022, PEVAP). N.J.M., D.E.S. and M.Z. acknowledge support from a UKRI Future Leaders Fellowship [Grant MR/T040866/1], a Science and Technology Facilities Funding Council Nucleus Award [Grant 954 ST/T000082/1], and the Leverhulme Trust through a research project grant [RPG-2020-82]. P.J.W. acknowledges support from the UK Science and Technology Facilities Council (STFC) through consolidated grants ST/T000406/1 and ST/X001121/1. V.P. acknowledges support from the UKRI Future Leaders Fellowship grant MR/S035214/1 and UKRI Science and Technology Facilities Council (STFC) through the consolidated grant ST/X001121/1. C.E.F. acknowledges support from the Swiss National Science Foundation (SNSF) Mobility Fellowship [grant no. P500PT_203110]. E.A. would like to thank C. McDonald and D. Veras for insightful discussions.
@@
The PLATON atmosphere models and PandExo spectral uncertainties will be made available on Zenodo.

### v2

@@ Abstract
A primary objective of exoplanet atmosphere characterisation is to learn about planet formation and evolution, however, this is challenged by degeneracies. To determine whether differences in atmospheric composition can be reliably traced to differences in evolution, we are undertaking a transmission spectroscopy survey with JWST to compare the compositions of a sample of hot Jupiters that have different orbital alignments around F stars above the Kraft break. Under the assumption that aligned planets migrate through the inner disc, while misaligned planets migrate after disc dispersal, the act of migrating through the inner disc should cause a measurable difference in the C/O between aligned and misaligned planets. We expect the amplitude and sign of this difference to depend on the amount of planetesimal accretion and whether silicates accreted from the inner disc release their oxygen. Here, we identify all known exoplanets that are suitable for testing this hypothesis, describe our JWST survey, and use noise simulations and atmospheric retrievals to estimate our survey’s sensitivity. With the selected sample of four aligned and four misaligned hot Jupiters, we will be sensitive to the predicted differences in C/O between aligned and misaligned hot Jupiters for a wide range of model scenarios.
@@
Instrumentation – Data methods – JWST – Exoplanets – Planet formation – Planet migration1 Introduction
@@
Penzlin & Booth et al. 2024 used simulations to demonstrate that the unconstrained values of key disc and planet formation parameters such as dust-to-gas mass, disc temperature, and the relative drift of dust to gas (Stokes-to-alpha number), create a degeneracy between C/O and [O/H] (which we refer to as metallicity, Z, throughout this paper) over a wide dynamic range. These uncertainties are hard to constrain robustly through independent observations. Therefore, we propose the best way to determine whether differences in formation history lead to a measurable difference in atmospheric C/O and metallicity is by comparing populations of planets for which we have independent evidence that they underwent different evolutionary pathways. Specifically, as Penzlin & Booth et al. 2024 show, planets that have undergone disc-free (high eccentricity) migration should have different C/O and metallicity to planets that have undergone disc migration. The idea behind this is that disc-migrated planets will accrete solids from the inner disc during their migration while disc-free, high-eccentricity migrated planets will not since they complete their migration after disc dispersal (Figure 1).
@@ 2 Target selection
To test whether atmospheric composition depends on migration history, we needed to construct a sample of planets with large predicted atmospheric signals and precisely measured obliquities. To do this, we drew from the TEPCAT catalogue111https://www.astro.keele.ac.uk/jkt/tepcat/obliquity.html, accessed 23rd May 2024 (Southworth, 2011) of 180 exoplanets with measured sky-projected obliquities, , primarily measured via the Rossiter-McLaughlin effect (Rossiter, 1924; McLaughlin, 1924). We define misaligned exoplanets as those with sky-projected obliquities ∘ and aligned planets as those with ∘ and only considered those planets with precisely measured obliquities (() ∘). While we based our sample selection on sky-projected obliquities, it is possible that planets with low sky-projected obliquities could, in fact, be misaligned in 3D space once the stellar inclination is resolved. However, of the 20 planets with measured sky-projected and 3D obliquities from the TEPCAT catalogue (Southworth, 2011), only one (5 % of the total) is aligned in 2D, sky-projected space and is misaligned in 3D space. This is encouraging for our target selection, which is based on sky-projected obliquities. Nevertheless, our sky-projected aligned sample could include planets that are misaligned in 3D space.
@@
Of these planets, we are acquiring new JWST observations of TrES-4b, KELT-7b, NGTS-2b, HAT-P-30b and WASP-15b through Programme GO 3838 (PIs: Kirk & Ahrer) with the other targets already observed or due to be observed by other approved programmes (WASP-17b, GTO 1353, PI: Lewis; HD 149026b, GTO 1274, PI: Lunine, Bean et al. 2023; WASP-94Ab, GO 3154, PI: Ahrer). All of these planets have eccentricities consistent with zero. With this combination of archival and new data, we will obtain (i) a four planet sample of aligned planets that likely migrated through the disc and (ii) a four planet sample of misaligned planets that ended their formation at orbits near and beyond the water ice line, and reach their final orbits after disc dispersal.
We note that our choice to divide aligned and misaligned planets at ∘ was motivated by Spalding & Winn (2022)’s definition. If we instead chose a more conservative definition of aligned planets, those with ∘, then HAT-P-41b ( ∘), XO-6b ( ∘) and WASP-3b ( ∘) would be classified as misaligned planets. Since these are not included in our observational sample, our observations are not sensitive to this definition of alignment.
A caveat to our experiment is the assumption that aligned planets are the result of disc migration and misaligned planets are the result of disc-free migration. However, there are other migration scenarios that could result in these obliquities. For example, coplanar high-eccentricity migration (Petrovich, 2015) can occur in planetary systems with two or more gas giants. In this case, both planets share a low mutual inclination that remains aligned with the stellar spin axis. The inner planet can become an aligned, circularised hot Jupiter while the outer planet is typically 1–3 times more massive and remains on a moderately eccentric orbit (). Zink & Howard (2023) used the California Legacy Survey to show that massive companions to hot Jupiters are ubiquitous at a population level. Furthermore, the eccentricity and mass distributions of these outer companions are compatible with the coplanar high-eccentricity migration scenario. While none of our aligned planets have a detected outer companion, outer companions could exist below the detection threshold. Separately, an inclined protoplanetary disc could result in misaligned hot Jupiters that underwent disc migration (e.g., Batygin, 2012; Spalding & Batygin, 2015). We discuss the possible implications of these alternative scenarios in Section 8.
@@ 3 Observing strategy
We will observe our targets in transmission using the JWST NIRSpec G395H instrument mode, which covers several absorption bands from the primary O and C bearing species in hot Jupiters (H2O, CO2 and CO, Figure 3 and e.g., Alderson et al. 2023). This approach will allow us to measure C/O and O/H and thus address whether these quantities differ between aligned and misaligned hot Jupiters. While NIRSpec/PRISM would cover the same molecular features in addition to covering bluer wavelengths (albeit at lower spectral resolution), it saturates more quickly than NIRSpec/G395H. The JWST Early Release Science observations of WASP-39b (K mag = 10.2) with NIRSpec/PRISM saturated at near-IR wavelengths (Rustamkulov et al., 2023), with the data synthesis analysis recommending the avoidance of partial saturation with PRISM (Carter et al., 2024). Since four of our five GO 3838 targets are brighter than WASP-39 (Table 1), we selected G395H to avoid saturation and to ensure a homogeneous data set across the sample. Similarly, the observations of WASP-17b and WASP-94Ab also make use of the G395H instrument. However, HD 149026b was observed with NIRCam in emission (Bean et al., 2023), and therefore, we will need to take this into account when we combine the full sample. We discuss the results of Bean et al. (2023) in Section 6.
@@ 4 Predictions from disc models and 1D atmosphere models
For our actual sample size of eight planets, we estimate a precision on C/O of (aligned planets) and (misaligned planets) for the solar metallicity, fiducial case (disc scenario 1). This would make the predicted difference between the aligned and misaligned planets of particularly challenging to measure (). However, if the planets are more metal-rich, similar to other planets in the literature, then we would be sensitive to the difference between the C/O of aligned and misaligned planets in the solar metallicity fiducial setup. Specifically, we estimate a precision on C/O of (aligned planets) and (misaligned planets) in this case, which would allow us to measure the predicted difference between the C/O of the two populations of at .
@@ 7 Additional science enabled by our homogeneous sample
In addition to the 1D forward and retrieval modelling we will perform for our sample, we will investigate the atmospheres of our exoplanets in 3D. Since all of our targets are on short orbital periods they are expected to be tidally locked. This leads to a large day-to-night temperature gradient between their permanent daysides and permanent nightsides which, in turn, drives a super-rotating equatorial jet that carries hot dayside gas to the nightside and cool nightside gas to the dayside (e.g., Showman & Polvani, 2010). The result is hotter evening limbs and cooler morning limbs, and hence different gas phase chemistry and cloud coverage at each limb. Using a combination of ingress/egress light curve fits and light curve fits with two planetary radii for the morning and evening limb, Espinoza et al. (2024) have measured limb differences in a planet’s transmission spectrum with JWST. We will apply the same techniques, using Tiberius and catwoman101010https://github.com/KathrynJones1/catwoman (Jones & Espinoza, 2020) to search for limb asymmetries in our data.
@@
Our investigations of 3D effects and photochemistry will be backed up by a comparison of each planet’s observed spectrum to the spectrum predicted by a 3D climate model (also known as a general circulation model, or GCM) of a hot Jupiter atmosphere. This comparison will reveal how the interactions between atmospheric circulation, radiative transfer, disequilibrium thermochemistry (e.g., Zamyatina et al., 2024) and photochemistry, in 3D, impact the exoplanets’ morning and evening limb spectra. Furthermore, this will allow us to determine to what extent the limb-averaged C/O and Z we measure from our transit spectra are representative of the overall atmosphere composition. Among the GCM outputs will be longitude-latitude maps of molecular abundances as a function of pressure and temperature, which we will use to inform our 1D models. The BOWIE-ALIGN programme sample could also serve as a high-quality observational reference for model intercomparison projects, such as those within the CUISINES framework (Fauchez et al., 2021; Sohl et al., 2024). The homogeneity of data would allow to benchmark single-column models and 3D GCMs in terms of e.g. their simulated transmission spectra, for aligned and misaligned hot Jupiters. This will help to validate our theoretical predictions and make the conclusions of the programme more robust.
@@ 8 Summary
: this result would indicate either that the majority of hot Jupiter assembly is completed before migration or that obliquity is not a sensitive tracer of migration for planets around F-type stars.
In this paper, we have described our observational test of these predictions, including the target selection and sample size, which we investigated using a combination of 256 1D atmospheric forward models, JWST noise simulations and atmospheric retrievals.
Our simulated transmission spectra show we will measure the predicted difference in C/O of aligned versus misaligned planets with confidences between 0 and , depending on where our targeted exoplanets sit in C/O–[O/H] parameter space (Figure 4). We will be sensitive to result (i) if the atmospheric metallicities approach solar, while the greater differences in C/O between the two populations for result (ii) means we will be sensitive to these over a wide range of metallicities. If we measure either result (i) or (ii), this would represent the first observational confirmation that atmospheric composition traces formation environment. If we observe super-solar metallicities, this would indicate the accretion of solid material, like the solar system, while sub-solar metallicities would suggest that hot Jupiter assembly differs from the giant planets in the solar system.
A caveat to our experiment is the assumption that disc-migration results in aligned systems while disc-free migration results in misaligned systems. Other scenarios, like coplanar high-eccentricity migration (Petrovich, 2015), could contaminate this distinction. If we measure a negligible difference between the compositions of aligned and misaligned planets, result (iii), this could be evidence for a common migration history and thus that obliquities are not a sensitive tracer of migration, even around F stars.
Ultimately, our comparative survey of a population of carefully-selected hot Jupiters will provide a rich dataset with which to test the dependence of atmospheric composition on formation and migration.
@@ Acknowledgements
J.K. acknowledges financial support from Imperial College London through an Imperial College Research Fellowship grant. J.E.O. and R.A.B are supported by Royal Society University Research Fellowships and A.B.T.P. received support from their Enhancement Awards. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Framework Programme (grant agreement no. 853022, PEVAP). N.J.M., D.E.S. and M.Z. acknowledge support from a UKRI Future Leaders Fellowship [Grant MR/T040866/1], a Science and Technology Facilities Council Nucleus Award [Grant 954 ST/T000082/1], and the Leverhulme Trust through a research project grant [RPG-2020-82]. P.J.W. acknowledges support from the UK Science and Technology Facilities Council (STFC) through consolidated grants ST/T000406/1 and ST/X001121/1. V.P. acknowledges support from the UKRI Future Leaders Fellowship grant MR/S035214/1 and UKRI Science and Technology Facilities Council (STFC) through the consolidated grant ST/X001121/1. C.E.F. acknowledges support from the Swiss National Science Foundation (SNSF) Mobility Fellowship [grant no. P500PT_203110] and the ERC under the European Union’s Horizon 2020 research and innovation program [grant agreement no 805445]. E.A. would like to thank C. McDonald and D. Veras for insightful discussions.
@@
The PLATON atmosphere models, retrieval results and PandExo spectral uncertainties are available on Zenodo at this link: https://zenodo.org/records/13902706.